Valenciaian:

The complex solution to a quadratic equation is x=-10± root -700/2. Write this solution in standard form, x=a±bi, where a and b are real numbers. Justify your answer by showing your work. HINT: Simplify the radical first.

jhonyy9:

\[\sqrt{\frac{ -700 }{ 2 }} = ?\]

Valenciaian:

??

jhonyy9:

\[\sqrt{\frac{ -7*100 }{ 2 }} = ?\]

jhonyy9:

\[\sqrt{100} = ?\]

jhonyy9:

\[\frac{ 7 }{ 2 } = ?\]

jhonyy9:

\[\sqrt{-1} = ?\]